OpenStudy (anonymous):

The area od a rectangle wall of a barn is 133 square feet. Its length is 12 feet longer that the width. Find the length and width of the wall of the barn.

OpenStudy (anonymous):

width = w length = w + 12 area = l * w = w(w + 12) = w^2 + 12w = 133 w^2 + 12w - 133 = 0 (w + 19)(w - 7) = 0 w = 7 (negative width makes no sense) l = 19

OpenStudy (anonymous):

thanks :D